OpenStudy (anonymous):

Solve \frac{ t }{5 } + 3 = 3.

OpenStudy (anonymous):

minus the 3 then multiply the 5

OpenStudy (anonymous):

\[\frac{ t }{ 5 } + 3 = 3 ?\]

OpenStudy (anonymous):

this should be an easy one

OpenStudy (anonymous):

yessss

OpenStudy (anonymous):

you dont get 6 you get 0 xD @Mateaus

OpenStudy (anonymous):

What the.

OpenStudy (anonymous):

t/5 +3 = 3 t/5 = 3 - 3 t/5 = 0 t = 0 * 5 t = 0

OpenStudy (anonymous):

t/5+3=3 t/5=0 t=0

OpenStudy (anonymous):

Sorry, I haven't sleep in one day yet. Is funny how the illusions of numbers work though lol

OpenStudy (anonymous):

that's it.? O.o

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yep